A new virus is released on the internet; the administrator of a department's Local Area Network (LAN) is given five minutes by a manager to estimate the impact. The administrator samples 12 of the PCs connected to the LAN and finds that 7 are infected; use proportion to estimate the number of infected PCs if there are a total of 117 PCs connected to the LAN.
Answer by stanbon(75887)

The administrator samples 12 of the PCs connected to the LAN and finds that 7 are infected; use proportion to estimate the number of infected PCs if there are a total of 117 PCs connected to the LAN
x/117 = 7/12
x = (7/12)*117
x = 68.25
Rounded down the estimated number of infected PCs is 68
-----------------
Cheers,
Stan H.